A company is developing a new high performance wax for cross country ski racing. In order to justify the price marketing​ wants, the wax needs to be very fast.​ Specifically, the mean time to finish their standard test course should be less than 55 seconds for a former Olympic champion. To test​ it, the champion will ski the course 8 times. The​ champion's times​ (selected at​ random) are 57.3​,63.1​, 46.1​, 50.4​,46.4​, 47.2​, 53.1​, and 42.3 seconds to complete the test course. Should they market the​ wax? Assume the assumptions and conditions for appropriate hypothesis testing are met for the sample. Use 0.05 as the​ P-value cutoff level.

t= ___

p-value=___
Answer #1

We normally assume that that champions times are distributed. high Let u= Population mean of new performance waxs timer. Wetnt = t8-1= ty P-value = p (tz *-7676). Using MS-Excel, p(t73-1.7676) - T.DLST (-7876, 7,1) 0.06 & x= 0.05 p-value (0.06) > a